OpenStudy (anonymous):

find the slope using points: (2, 2) and (-5, 4)

terenzreignz (terenzreignz):

You know, slope is all about differences ^_^ In fact, it is equal to the difference of the y-coordinates divided by the difference of the x-coordinates...

OpenStudy (anonymous):

\[m = \left(\begin{matrix}y ^{2}-y ^{1} \\ x ^{2}-x ^{1}\end{matrix}\right)\]

OpenStudy (anonymous):

substitute in and solve :)

terenzreignz (terenzreignz):

... -_- \[\Large m = \frac{y_2 - y_1}{x_2-x_1}\]